# Arithmetic aspects of discrete periodic Toda flows

### v2

@@ 3 How to recover the periodic box-ball flow
3.1 Periodic box-ball flow
@@
$\mathcal{B}:\mathcal{V}_{N}\longrightarrow\mathcal{V}_{N}$ (3.1)
explained in the Introduction, where the state space is given by $\mathcal{V}_{N}=\{y\in\{0,1\}^{N}|\sum_{i=1}^{N}y_{i}<N/2\}$. Very surprisingly, as shown in [14], the periodic box-ball flow can naturally be used to formulate the famous Riemann hypothesis.
@@
3.2 Tropical periodic Toda flow
@@
$\mathcal{T}^{\mathrm{trop}}:\mathbb{R}^{2n}\dashrightarrow\mathbb{R}^{2n},$ (3.2)
$(\mathbb{J}^{t},\mathbb{W}^{t})=(J_{1}^{t},\dots,J_{n}^{t},W_{1}^{t},\dots,W_{n}^{t})\longmapsto(\mathbb{J}^{t+1},\mathbb{W}^{t+1}),$ (3.3)
@@
$\begin{split}J^{t+1}_{i}&=\mathrm{min}\{W^{t}_{i},J_{i}^{t}+X_{i}^{t}\},\\ W^{t+1}_{i}&=J^{t}_{i+1}+W^{t}_{i}-J^{t+1}_{i},\end{split}$ (3.4)
@@
3.3 From periodic box-ball to tropical periodic Toda
@@
$\eta:\mathcal{V}_{N}\longrightarrow\bigsqcup_{n=1}^{\lfloor N/2\rfloor}\mathbb{N}^{2n}$ (3.7)
from the phase space of the periodic box-ball system to the disjoint union of the phase spaces of the tropical periodic Toda system. It is important to note that the subset
$\mathcal{V}_{N,n}=\{y\in\mathcal{V}_{N}\ |\ \eta(y)\in\mathbb{N}^{2n}\}$ (3.8)
@@
Let us explain the definition of $\eta$ with an example. Consider the state $y_{13}=1100011000101\in\mathcal{V}_{13}$ from the Introduction. First of all $n=4$ is given by the number of connected components of $1$’s. Then we record (from left to right) the sizes of the connected components of $1$’s and 00’s, leading to $\mathbb{J}_{y_{13}}=(2,2,1,1)$ and $\mathbb{Q}_{y_{13}}=(3,3,1,0)$.
@@
${\mathcal{V}_{N,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$$\scriptstyle{\mathcal{B}}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{T}^{\mathrm{trop}}}$${\mathcal{V}_{N,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}}$ (3.10)
@@
$\eta:\mathcal{V}_{N,n}\longrightarrow\mathbb{N}^{2n}{/{\langle\sigma\rangle}}$ (3.11)
@@
3.4 Lifting tropical Toda to discrete Toda
@@
$I_{\epsilon,i}=e^{-J_{i}/\epsilon}\text{ and }V_{\epsilon,i}=e^{-W_{i}/\epsilon}.$ (3.12)
@@
$\lim_{\epsilon\to 0}-\epsilon\,\mathrm{log}(I_{\epsilon,i})=J_{i}\text{ and }\lim_{\epsilon\to 0}-\epsilon\,\mathrm{log}(V_{\epsilon,i})=W_{i}.$ (3.13)
@@
${\mathbb{N}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{\epsilon}}$$\scriptstyle{\mathfrak{T}^{\mathrm{trop}}}$${\mathbb{R}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{T}}$${\mathbb{N}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{\epsilon}}$${\mathbb{R}^{2n}.}$ (3.14)
Instead, we prefer to consider an algebraic lift. Introduce the variable $T$ and define
$\mathfrak{I}_{T}:(\mathbb{J},\mathbb{W})\in\mathbb{N}^{2n}\longmapsto(\mathbb{I}_{T},\mathbb{V}_{T})\in\mathbb{Q}(T)^{2n}$ (3.15)
@@
$I_{T,i}=T^{J_{i}}\text{ and }V_{T,i}=T^{W_{i}}.$ (3.16)
In this set-up the tropicalization procedure corresponds to the $T$-adic valuation $\nu_{T}$ on $\mathbb{Q}(T)$, i.e., we recover $(\mathbb{J},\mathbb{W})$ from $(\mathbb{I}_{T},\mathbb{V}_{T})$ simply by
$J_{i}=\nu_{T}(I_{T,i})\text{ and }W_{i}=\nu_{T}(W_{T,i}).$ (3.17)
In particular this means the the discrete periodic Toda flow over $\mathbb{Q}(T)$ contains the tropical periodic Toda flow via
@@
3.5 All together
Let us define $\mathcal{V}_{N,\mathcal{C}}=\eta^{-1}(\nu_{T}(\mathcal{R}_{\mathcal{C}}))$. Summarising everything said so far, we obtain the following description of the periodic box-ball flow in terms Mumford’s $\mathrm{Jac(\mathcal{C})}$ and the Gauß composition law.
@@ 3.5 All together
${\mathcal{V}_{N,\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$$\scriptstyle{\mathcal{B}}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{T}}$$\scriptstyle{\mathcal{T}^{\mathrm{trop}}}$${\mathbb{Q}(T)^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{\mathcal{C}}}$$\scriptstyle{\mathcal{T}}$${\mathrm{Jac}(\mathcal{C})/\langle\mathfrak{pa}\rangle\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot\boxplus\mathfrak{T}}$${\mathcal{V}_{N,\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{T}}$${\mathbb{Q}(T)^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{\mathcal{C}}}$${\mathrm{Jac}(\mathcal{C})/\langle\mathfrak{pa}\rangle}$ (3.19)
@@
3.6 Interlude: Tropical Mumford Jacobian
We continue to work with the lift $\mathfrak{I}_{T}:\mathbb{N}^{2n}\longrightarrow\mathbb{Q}(T)^{2n}$. For background on tropical polynomials we refer to [13].
Definition 3.1.
Given $F(x)=\sum_{i=0}^{n}f_{i}x^{i}\in\mathbb{Q}(T)[x]$, we define its tropical polynomial by
$F^{\mathrm{trop}}(x)=\mathrm{min}\{\nu_{T}(a_{0}),\nu_{T}(a_{1})+x,\dots,\nu_{T}(a_{n})+nx\}.$ (3.20)
@@ 3.6 Interlude: Tropical Mumford Jacobian
$\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}=\{([u^{\mathrm{trop}},v^{\mathrm{trop}}],n)\ |\ ([u,v],n)\in\mathrm{Jac(\mathcal{C})}\}$ (3.21)
@@
$\mathcal{R}_{\mathcal{C}}^{\mathrm{trop}}=\{\nu_{T}((\mathbb{I},\mathbb{V}))|(\mathbb{I},\mathbb{V})\in\mathcal{R}_{\mathcal{C}}\}\subset\mathbb{N}^{2n}.$ (3.22)
Proposition 3.3 (Tropical inverse Abel-Jacobi map).
@@
$(\Psi^{-1}_{\mathcal{C}})^{\mathrm{trop}}:\mathrm{Im}(\Psi_{\mathcal{C}})^{\mathrm{trop}}\longrightarrow\mathcal{R}_{\mathcal{C}}^{\mathrm{trop}}$ (3.23)
@@
Proposition 3.4 (Linearization of tropical Toda flow).
@@
$\boxplus\mathfrak{T}^{\mathrm{trop}}:\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}\longrightarrow\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}$ (3.25)
@@
${\mathrm{Jac(\mathcal{C})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot\boxplus\mathfrak{T}}$${\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\boxplus\mathfrak{T}^{\mathrm{trop}}}$${\mathrm{Jac(\mathcal{C})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}.}$ (3.26)
@@
For this we exploit again Lemma 2.5. We’re just using transport of structure at this point. ∎
Remark 3.6.
It should be interesting to investigate the Gauß composition law and the theory of division polynomials in the framework of tropical curves and Jacobians. We do not pursuit this direction in the present paper.
4 Integrality result and applications
Theorem 4.1.
The discrete periodic Toda flow induces a flow
$\mathcal{T}:\mathbb{Z}[T]_{(T)}^{2n}\dashrightarrow\mathbb{Z}[T]_{(T)}^{2n},$ (4.1)
which recovers the periodic box-ball flow using the $T$-adic valuation $\nu_{T}$. Here $\mathbb{Z}[T]_{(T)}$ denotes the localization of $\mathbb{Z}[T]$ at the ideal $(T)$ which is a local ring [^8] In particular we have $T^{-1}\notin\mathbb{Z}[T]_{(T)}$..
Proof.
For $(\mathbb{I}^{t},\mathbb{V}^{t})\in\mathbb{Z}[T]_{(T)}^{2n}$ we need to show that
$\nu_{T}(I^{t+1}_{i}),\nu_{T}(V^{t+1}_{i})\geq 0$ (4.2)
$\nu_{T}(I_{o}),\nu_{T}(V_{o})\geq\nu_{T}(\sigma^{k}(u^{t+1}_{x}))$ (4.3)
for all $k\in\{1,\dots,n\}$. ∎
Corollary 4.1.
The substitution $T\mapsto p$, for $p$ a (big enough) prime number, leads to the flow
$\mathcal{T}_{p}:\mathbb{Z}_{(p)}^{2n}\dashrightarrow\mathbb{Z}_{(p)}^{2n},$ (4.4)
$\mathcal{T}_{p}:\mathbb{Z}_{p}^{2n}\dashrightarrow\mathbb{Z}_{p}^{2n}.$ (4.5)
Remark 4.1.
5 Outlook
p-adic discrete periodic Toda flow
Berkovich spaces over $\mathbb{Z}$
Let us finish our paper with the following somewhat speculative
Question 5.1.
Appendix A Monodromy matrix approach
$M_{n}=\begin{bmatrix}m^{1}_{n+1}&m^{1}_{n}\\ m^{2}_{n+1}&m^{2}_{n}\end{bmatrix},$ (A.1)
$m_{j}=(I_{j+1}+V_{j}+x)m_{j-1}-I_{j}V_{j}m_{j-2}$ (A.2)
together with the two initial conditions
$(m^{1}_{1},m^{1}_{0})=(1,0)\text{ and }(m^{2}_{1},m^{2}_{0})=(0,1),$ (A.3)
leading to the two relations
$h_{n}(x)=\mathrm{tr}(M_{n})=m^{1}_{n+1}+m^{2}_{n}$ (A.4)
and
$f_{n}=|M_{n}|=m^{1}_{n+1}m^{2}_{n}-m^{1}_{n}m^{2}_{n+1}.$ (A.5)
The connection to our set-up is given by the next
Lemma A.1.
For $n\geq 2$, we have
$(m^{1}_{n},m^{2}_{n})$ $=(\sigma^{2-n}(u_{n}),\sigma^{2-n}(-v_{n})),$ (A.6)
$(m^{1}_{n+1},m^{2}_{n+1})$ $=(\sigma^{2-n}(h_{n}+v_{n}),\sigma^{2-n}(-I_{n}V_{n}\sigma(u_{n}))).$ (A.7)
Proof.
We leave this as an exercise. ∎

### v3

@@ 3. How to recover the periodic box-ball flow
3.1. Periodic box-ball flow
@@
$\mathcal{B}:\mathcal{V}_{N}\longrightarrow\mathcal{V}_{N}$
explained in the Introduction, where the state space is given by $\mathcal{V}_{N}=\{y\in\{0,1\}^{N}|\sum_{i=1}^{N}y_{i}<N/2\}$. Remarkably, as shown in [14], the periodic box-ball flow can naturally be used to formulate the famous Riemann hypothesis.
@@
3.2. Tropical periodic Toda flow
@@
$\mathcal{T}^{\mathrm{trop}}:\mathbb{R}^{2n}\dashrightarrow\mathbb{R}^{2n},$
$(\mathbb{J}^{t},\mathbb{W}^{t})=(J_{1}^{t},\dots,J_{n}^{t},W_{1}^{t},\dots,W_{n}^{t})\longmapsto(\mathbb{J}^{t+1},\mathbb{W}^{t+1}),$
@@
(3.1) $\begin{split}J^{t+1}_{i}&=\mathrm{min}\{W^{t}_{i},J_{i}^{t}+X_{i}^{t}\},\\ W^{t+1}_{i}&=J^{t}_{i+1}+W^{t}_{i}-J^{t+1}_{i},\end{split}$
@@
3.3. From periodic box-ball to tropical periodic Toda
@@
$\eta:\mathcal{V}_{N}\longrightarrow\bigsqcup_{n=1}^{\lfloor N/2\rfloor}\mathbb{N}^{2n}$
from the phase space of the periodic box-ball system to the disjoint union of the phase spaces of the tropical periodic Toda system. Importantly, the subset
$\mathcal{V}_{N,n}=\{y\in\mathcal{V}_{N}\ |\ \eta(y)\in\mathbb{N}^{2n}\}$
@@
We explain the definition of $\eta$ with an example. Consider the state $y_{13}=1100011000101\in\mathcal{V}_{13}$ from the Introduction. First of all $n=4$ is given by the number of connected components of $1$’s. Then we record (from left to right) the sizes of the connected components of $1$’s and 0’s, leading to $\mathbb{J}_{y_{13}}=(2,2,1,1)$ and $\mathbb{Q}_{y_{13}}=(3,3,1,0)$.
@@
${\mathcal{V}_{N,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}\scriptstyle{\mathcal{B}}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{T}^{\mathrm{trop}}}$${\mathcal{V}_{N,n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}}$
@@
$\eta:\mathcal{V}_{N,n}\longrightarrow\mathbb{N}^{2n}{/{\langle\sigma\rangle}}$
@@
3.4. Lifting tropical Toda to discrete Toda
@@
$I_{\epsilon,i}=e^{-J_{i}/\epsilon}\text{ and }V_{\epsilon,i}=e^{-W_{i}/\epsilon}.$
@@
$\lim_{\epsilon\to 0}-\epsilon\,\mathrm{log}(I_{\epsilon,i})=J_{i}\text{ and }\lim_{\epsilon\to 0}-\epsilon\,\mathrm{log}(V_{\epsilon,i})=W_{i}.$
@@
${\mathbb{N}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{\epsilon}}$$\scriptstyle{\mathcal{T}^{\mathrm{trop}}}$${\mathbb{R}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{T}}$${\mathbb{N}^{2n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{\epsilon}}$${\mathbb{R}^{2n}.}$
Instead, we prefer to consider an algebraic lift. Introduce the variable $q$ and define
(3.2) $\mathfrak{I}_{q}:(\mathbb{J},\mathbb{W})\in\mathbb{N}^{2n}\longmapsto(\mathbb{I}_{q},\mathbb{V}_{q})\in\mathbb{Q}(q)^{2n}$
@@
$I_{q,i}=q^{J_{i}}\text{ and }V_{q,i}=q^{W_{i}}.$
In this set-up the tropicalization procedure corresponds to the $q$-adic valuation $\nu_{q}$ on $\mathbb{Q}(q)$, i.e., we recover $(\mathbb{J},\mathbb{W})$ from $(\mathbb{I}_{q},\mathbb{V}_{q})$ simply by
$J_{i}=\nu_{q}(I_{q,i})\text{ and }W_{i}=\nu_{q}(V_{q,i}).$
In particular this means the discrete periodic Toda flow over $\mathbb{Q}(q)$ contains the tropical periodic Toda flow via
@@
3.5. All together
We define $\mathcal{V}_{N,\mathcal{C}}=\eta^{-1}(\nu_{q}(\mathcal{R}_{\mathcal{C}}))$. Summarising everything said so far, we obtain the following description of the periodic box-ball flow in terms Mumford’s $\mathrm{Jac(\mathcal{C})}$ and the Gauß composition law.
@@ 3.5. All together
${\mathcal{V}_{N,\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}\scriptstyle{\mathcal{B}}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{q}}$$\scriptstyle{\mathcal{T}^{\mathrm{trop}}}$${\mathbb{Q}(q)^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{\mathcal{C}}}$$\scriptstyle{\mathcal{T}}$${\mathrm{Jac}(\mathcal{C})/\langle\mathfrak{pa}\rangle\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot\boxplus\mathfrak{T}}$${\mathcal{V}_{N,\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta}$${\mathbb{N}^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{I}_{q}}$${\mathbb{Q}(q)^{2n}{/{\langle\sigma\rangle}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi_{\mathcal{C}}}$${\mathrm{Jac}(\mathcal{C})/\langle\mathfrak{pa}\rangle}$
@@
4. Integrality result and applications
Theorem 4.1.
The discrete periodic Toda flow induces a flow
$\mathcal{T}_{\mathrm{integral}}:\mathbb{Z}[q]_{(q)}^{2n}\dashrightarrow\mathbb{Z}[q]_{(q)}^{2n},$
which recovers the periodic box-ball flow using the $q$-adic valuation $\nu_{q}$. Here $\mathbb{Z}[q]_{(q)}$ denotes the localization of $\mathbb{Z}[q]$ at the ideal $(q)$, which is a local ring [^7] In particular we have $q^{-1}\notin\mathbb{Z}[q]_{(q)}$..
Proof.
For $(\mathbb{I}^{t},\mathbb{V}^{t})\in\mathbb{Z}[q]_{(q)}^{2n}$ we need to show that
$\nu_{q}(I^{t+1}_{i}),\nu_{q}(V^{t+1}_{i})\geq 0$
$\nu_{q}(I_{\Pi}),\nu_{q}(V_{\Pi})\geq\nu_{q}(\sigma^{k}(u^{t+1}_{x}))$
for all $k\in\{1,\dots,n\}$. ∎
Corollary 4.1.
The substitution $q\mapsto p$, for $p$ a (big enough) prime number, leads to the flow
$\mathcal{T}_{p}:\mathbb{Z}_{(p)}^{2n}\dashrightarrow\mathbb{Z}_{(p)}^{2n},$
$\mathcal{T}_{p}:\mathbb{Z}_{p}^{2n}\dashrightarrow\mathbb{Z}_{p}^{2n}.$
Remark 4.1.
5. Outlook
p-adic discrete periodic Toda flow
Berkovich spaces over $\mathbb{Z}$
We finish our paper with the following somewhat speculative
Question 5.1.
Appendix A Monodromy matrix approach
$M_{n}=\begin{bmatrix}m^{1}_{n+1}&m^{1}_{n}\\ m^{2}_{n+1}&m^{2}_{n}\end{bmatrix}$,
$m_{j}=(I_{j+1}+V_{j}+x)m_{j-1}-I_{j}V_{j}m_{j-2}$
together with the two initial conditions
$(m^{1}_{1},m^{1}_{0})=(1,0)\text{ and }(m^{2}_{1},m^{2}_{0})=(0,1),$
leading to the two relations
$h_{n}(x)=\mathrm{tr}(M_{n})=m^{1}_{n+1}+m^{2}_{n}$
and
$f_{n}=|M_{n}|=m^{1}_{n+1}m^{2}_{n}-m^{1}_{n}m^{2}_{n+1}.$
The connection to our set-up is given by the next
Lemma A.1.
For $n\geq 2$, we have
$(m^{1}_{n},m^{2}_{n})$ $=(u_{n}^{\sigma^{2-n}},-v_{n}^{\sigma^{2-n}}),$
$(m^{1}_{n+1},m^{2}_{n+1})$ $=((h_{n}+v_{n})^{\sigma^{2-n}},(-I_{n}V_{n}u_{n}^{\sigma})^{\sigma^{2-n}}).$
Proof.
We leave this as an exercise. ∎
Appendix B Cantor’s division polynomials
We want to explain the following theorem due to Cantor
Theorem B.1 ([2] Theorem 8.7).
$r\cdot[X-x,y]=[\Delta_{r}(\tfrac{a-X}{4b^{2}},E_{r}(\tfrac{a-X}{4b^{2}})]\in\mathrm{Jac(\mathcal{D})}.$
B.1. Discrete open Toda flow
Set E(z)=F(az)E(z)=F(a-z) and consider the formal power series
$S(z)=\sum_{i\geq 0}s_{i}z^{i}=(-1)^{g+1}\sqrt{E(z)}.$
Define the Hankel determinants
$h_{m,n}=(-1)^{\binom{n}{2}}|(s_{m+i-j})_{i,j=1}^{n}|.$
(B.1) $h_{m-1,n}h_{m+1,n}=h_{m,n-1}h_{m,n+1}+h_{m,n}^{2}.$
Remark B.1.
The relation (B.1) is equivalent to the open discrete Toda equation in the $\tau$-function formalism, as explained in [7, 4].
B.2. Cantor’s division polynomials
$f_{r}=\left\{\begin{array}[]{cc}h_{m_{r},n_{r}}&,r\geq g\\ \ 0&,r<g\end{array}.\right$.
$f_{r-1}A^{i}_{r+1}(z)=f_{r}A^{i}_{r}(z)-zf_{r+1}A^{i}_{r-1}(z)$
with initial conditions
$(A^{1}_{g-1},A^{1}_{g})=(-z^{g-1},-z^{g})\text{ and }(A^{2}_{g-1},A^{2}_{g})=(0,0).$
$D_{r}(z)=-(A^{1}_{r}(z)^{2}-A^{2}_{r}(z)^{2}E(z))/z^{r}$
and
$\tilde{E}(z)=\tfrac{zf_{r-1}f_{r+1}}{2f_{r}^{2}}(\tfrac{D_{r+1}(z)}{f_{r+1}^{2}}-\tfrac{D_{r-1}(z)}{f_{r-1}^{2}})\text{ mod }D_{r}(z).$
$\Delta_{r}(z)=(2b)^{2\nu_{r}}D_{r}(4b^{2}z)\text{ and }E_{r}(z)=\tilde{E}_{r}(4b^{2}z),$
where $\nu_{r}=\binom{r}{2}-\binom{g}{2}$.
B.3. Back to our set-up
$t\cdot\mathfrak{T}\in\mathrm{Jac(\mathcal{C})}.$
$\mathfrak{T}_{\alpha}=[x_{\alpha}-\tfrac{1}{\alpha},\tfrac{V_{\Pi}-I_{\Pi}}{\beta(-\alpha)^{g+1}}]\in\mathrm{Jac}(\mathcal{C}_{\alpha})$
to which we can apply Cantor’s division polynomials (for $t>g$).
Remark B.2.
As far as the author knows, there is surprisingly no direct (algebraic) approach to division polynomials for real hyperelliptic curves in the literature. Interestingly, $p$-adic methods might pave the way to such a theory.
Appendix C Tropical Mumford Jacobian
We continue to work with the lift $\mathfrak{I}_{q}:\mathbb{N}^{2n}\longrightarrow\mathbb{Q}(q)^{2n}$ defined in (3.2). For background on tropical polynomials we refer to [13].
Definition C.1.
Given $F(x)=\sum_{i=0}^{n}f_{i}x^{i}\in\mathbb{Q}(q)[x]$, we define its tropical polynomial by
$F^{\mathrm{trop}}(x)=\mathrm{min}\{\nu_{q}(a_{0}),\nu_{q}(a_{1})+x,\dots,\nu_{q}(a_{n})+nx\}.$
@@ Appendix C Tropical Mumford Jacobian
$\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}=\{([u^{\mathrm{trop}},v^{\mathrm{trop}}],n)\ |\ ([u,v],n)\in\mathrm{Jac(\mathcal{C})}\}$
@@
$\mathcal{R}_{\mathcal{C}}^{\mathrm{trop}}=\{\nu_{q}((\mathbb{I},\mathbb{V}))|(\mathbb{I},\mathbb{V})\in\mathcal{R}_{\mathcal{C}}\}\subset\mathbb{N}^{2n}.$
Proposition C.1 (Tropical inverse Abel-Jacobi map).
@@
$(\Psi^{-1}_{\mathcal{C}})^{\mathrm{trop}}:\mathrm{Im}(\Psi_{\mathcal{C}})^{\mathrm{trop}}\longrightarrow\mathcal{R}_{\mathcal{C}}^{\mathrm{trop}}$
@@
Proposition C.2 (Linearization of tropical Toda flow).
@@
$\boxplus\mathfrak{T}^{\mathrm{trop}}:\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}\longrightarrow\mathrm{Jac(\mathcal{C})}^{\mathrm{trop}}$
@@
[…]
@@
For this we exploit again Lemma 2.5 and use transport of structure. ∎
Remark C.1.
It should be interesting to investigate the Gauß composition law and the theory of division polynomials in the framework of tropical curves and Jacobians. We do not pursue this direction in this paper.